# Central limit theorems for interacting innovation processes, related statistical tools and general results

### v3

@@ Appendix A General results
Hence, combining (see (Crimaldi et al., 2016, Lemma 4.1)) this limit with the limit relation
@@
A key point in the proof of Theorem A.3 is the convergence of the left-hand term of (A.7) toward a bounded limit, which is ensured by assumption (v), i.e. ${\mathcal{R}e}(\phi_{2}^{*})<\phi^{*}/{2}$. On the other hand, when ${\mathcal{R}e}(\phi_{2}^{*})\geq\phi^{*}/2$, this quantity can diverge to $\infty$, implying the divergence of $t^{\phi^{*}/2}(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})$. In fact, since $\Phi$ is a real matrix, in its spectrum we have $\phi_{2}^{*}$ and its conjugate, so that $F_{k+1,t}(x)F_{k+1,t}(y)=|F_{k+1,t}(x)|^{2}$ when we take $x=\alpha_{2}^{*}=\phi^{*}-\phi_{2}^{*}$ and for $y$ its conjugate (or $y=x$ if $x$ is real). Now, we observe that $|F_{k+1,t}(\alpha_{2}^{*})|^{2}$ increases with ${\mathcal{R}e}(\phi_{2}^{*})$ and the corresponding limit in (A.7), that is $1/(2{\mathcal{R}e}(\alpha_{2}^{*})-\phi^{*})=1/(\phi^{*}-2{\mathcal{R}e}(\phi_{2}^{*}))$, goes to $\infty$ for ${\mathcal{R}e}(\phi_{2}^{*})\uparrow\phi^{*}/2$. As a consequence of these two facts, we can conclude that the limit of the left-hand term of (A.7) is necessarily $\infty$ whenever ${\mathcal{R}e}(\phi_{2}^{*})\geq\phi^{*}/2$.
@@
When $N=1$, we simply have $(A_{t},B_{t})=(\widetilde{A}_{t},\widetilde{A}_{t}/\phi^{*}+o(t^{-\phi^{*}/2}))$ and so the statement follows from the clt for $\widetilde{A}_{t}$. When $N\geq 2$, we have
@@
Corollary A.5.
(CLT for $(\widetilde{B}_{t})$ when $N\geq 2$)
@@
Corollary A.6.
(CLT for $({\mathbf{A}}_{t}/({\mathbf{x}}^{\top}{\mathbf{A}}_{t}),{\mathbf{B}}_{t}/({\mathbf{x}}^{\top}{\mathbf{B}}_{t}))$ when $N\geq 2$)
@@
Indeed, we have $Q_{\mathbf{x}}{\mathbf{u}}{\mathbf{u}}^{\top}=0$ and so
@@ B.3. Hypothesis tests on the matrix $\Phi$ based on $({\mathbf{A}}_{t})$
(B.1) ${{\mathbf{T}}}_{t}=t^{\phi^{*}/2}g(\widetilde{A}_{t})^{-1/2}L^{11}(O^{11})^{\top}\,U\,V^{\top}\,{{\mathbf{A}}}_{t}=t^{\phi^{*}/2}g(\widetilde{A}_{t})^{-1/2}L^{11}(O^{11})^{\top}\,\widehat{{{\mathbf{A}}}}_{t}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},I_{N-1}\right)$
@@
In summary, whenever ${\mathbf{u}}_{0}={\mathbf{u}}_{1}$ we have under $H_{1}$ that ${{\mathbf{T}}}_{t}$ is still asymptotically normally distributed with covariance matrix $L^{11}(O^{11})^{\top}UV^{\top}{\mathcal{M}}_{det}^{11}VU^{\top}O^{11}(L^{11})^{\top}$, where ${\mathcal{M}}_{det}^{11}$ is referred to the eigen-structure of $\Phi_{1}$, while $L^{11}(O^{11})^{\top}UV^{\top}$ is referred to $\Phi_{0}$. Thus, the distance between the identity $I$ and $L^{11}(O^{11})^{\top}UV^{\top}{\mathcal{M}}_{det}^{11}VU^{\top}O^{11}(L^{11})^{\top}$ describes the relation between the asymptotic distribution of $\|{{\mathbf{T}}}_{t}\|^{2}$ under $H_{0}$ and the one under $H_{1}$, which determines the power of the test. For instance, note that $E[\|{{\mathbf{T}}}_{t}\|^{2}]=(N-1)$ under $H_{0}$, while $E[\|{{\mathbf{T}}}_{t}\|^{2}]$ is equal to the trace of $L^{11}(O^{11})^{\top}UV^{\top}{\mathcal{M}}_{det}^{11}VU^{\top}O^{11}(L^{11})^{\top}$ under $H_{1}$.
@@ B.4. Hypothesis tests on the matrix $\Phi$ based on $({\mathbf{B}}_{t})$
(B.2) ${{\mathbf{T}}}_{t}=t^{\phi^{*}/2}g(\phi^{*}\widetilde{B}_{t})^{-1/2}L^{33}(O^{33})^{\top}{\mathbf{B}}^{\prime}_{t}=t^{\phi^{*}/2}g(\phi^{*}\widetilde{B}_{t})^{-1/2}L^{33}(O^{33})^{\top}\,U\,V^{\top}\,{{\mathbf{B}}}_{t}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},I_{N-1}\right)$
@@ Appendix D Technical results
For the readear’s convenience, we here recall some results used in the above proofs. (For a synthetic review about the notion of stable convergence and its variants, see for instance the appendix of Aletti et al. (2020) and the references recalled therein.)
Theorem D.1.
(A consequence of (Crimaldi, 2009, Theorem A.1))
@@
Lemma D.2.
((Aletti et al., 2024, Supplementary material))
@@
Theorem D.3.
(A consequence of (Crimaldi and Pratelli, 2005, Proposition 3.1))
@@
Recall that, according to the notation used here, the above condition 2) means that $\sup_{m_{0}\leq k\leq t}|s_{t}{\mathbf{T}}_{t,k+1}|=s_{t}\sup_{m_{0}\leq k\leq t}\sum_{i=1}^{d}|T_{t,k+1,i}|\leq Cy_{t}$ for a suitable constant $C$ and $t$ large enough, so that we trivially have the convergence in $L^{1}$ of $\sup_{m_{0}\leq k\leq t}|s_{t}{\mathbf{T}}_{t,k+1}|$ toward zero, as requested in (Crimaldi and Pratelli, 2005, Proposition 3.1).
@@
Lemma D.4.
(A consequence of (Berti et al., 2011, Lemma 1))

### v4

@@ Appendix A General results
Hence, combining (see (Crimaldi et al.,, 2016, Lemma 4.1)) this limit with the limit relation
@@
Relaxing the assumption of diagonalizability of the matrix $\Phi$ would lead to analogous results. Indeed, it is enough to work with the associated Jordan matrix, instead of the matrix $D$ in (A.3). However, doing so would substantially increase the notational complexity and the length of the paper, making it unnecessarily cumbersome. Note that the fundamental ideas of the proof will remain the same, only the computations and the final formulas will become more complex and difficult to read. To get an idea, we refer to Yang et al., (2025), where, for a different context, the assumption of a diagonalizable matrix is replaced by the assumption of a block upper-triangular matrix. Note that the proofs are still based on (Aletti et al.,, 2017, Lemma A.4), (Aletti et al.,, 2019, Lemma A.5) and its analogous (Yang et al.,, 2025, Lemma A.5). Moreover, in the case of a reducible matrix $\Phi$, we refer to Aletti et al., (2023, 2024); Aletti and Ghiglietti, (2017).
Remark A.2.
A key point in the proof of Theorem A.3 is the convergence of the left-hand term of (A.7) toward a bounded limit, which is ensured by assumption (v), i.e. ${\mathcal{R}e}(\phi_{2}^{*})<\phi^{*}/{2}$. On the other hand, when ${\mathcal{R}e}(\phi_{2}^{*})\geq\phi^{*}/2$, this quantity can diverge to $\infty$, implying the divergence of $t^{\phi^{*}/2}(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})$. In fact, since $\Phi$ is a real matrix, in its spectrum we have $\phi_{2}^{*}$ and its conjugate, so that $F_{k+1,t}(x)F_{k+1,t}(y)=|F_{k+1,t}(x)|^{2}$ when we take $x=\alpha_{2}^{*}=\phi^{*}-\phi_{2}^{*}$ and for $y$ its conjugate (or $y=x$ if $x$ is real). Now, we observe that $|F_{k+1,t}(\alpha_{2}^{*})|^{2}$ increases with ${\mathcal{R}e}(\phi_{2}^{*})$ and the corresponding limit in (A.7), that is $1/(2{\mathcal{R}e}(\alpha_{2}^{*})-\phi^{*})=1/(\phi^{*}-2{\mathcal{R}e}(\phi_{2}^{*}))$, goes to $\infty$ for ${\mathcal{R}e}(\phi_{2}^{*})\uparrow\phi^{*}/2$. As a consequence of these two facts, we can conclude that the limit of the left-hand term of (A.7) is necessarily $\infty$ whenever ${\mathcal{R}e}(\phi_{2}^{*})\geq\phi^{*}/2$. This means that, in that case, the rate of convergence is slower than $t^{\phi^{*}/2}$.
@@
When $N=1$, we simply have $(A_{t},B_{t})=(\widetilde{A}_{t},\widetilde{A}_{t}/\phi^{*}+o(t^{-\phi^{*}/2}))$ and so the statement follows from Theorem A.1 and Theorem A.2 for $\widetilde{A}_{t}$. When $N\geq 2$, we have
@@
Corollary A.5 (CLT for $(\widetilde{B}_{t})$ when $N\geq 2$).
@@
Corollary A.6 (CLT for $({\mathbf{A}}_{t}/({\mathbf{x}}^{\top}{\mathbf{A}}_{t}),{\mathbf{B}}_{t}/({\mathbf{x}}^{\top}{\mathbf{B}}_{t}))$ when $N\geq 2$).
@@
Indeed, we have $Q_{\mathbf{x}}{\mathbf{u}}{\mathbf{u}}^{\top}=0$ and so, since ${\mathbf{x}}^{\top}{\mathbf{A}}_{t}\to({\mathbf{x}}^{\top}{\mathbf{u}})\widetilde{A}_{\infty}$ and ${\mathbf{x}}^{\top}{\mathbf{B}}_{t}\to({\mathbf{x}}^{\top}{\mathbf{u}})\widetilde{A}_{\infty}/\phi^{*}$, we have
@@ B.1. Confidence intervals for the random limit $\widetilde{A}_{\infty}$ based on $({\mathbf{A}}_{t})$
Since the convergence in Theorem A.1 is also in the sense of the almost sure conditional convergence with respect to $({\mathcal{F}}_{t})$, the above interval can be recovered as a credible interval (see Fortini and Petrone, (2020)) in a Bayesian framework.
@@ B.3. Hypothesis tests on the matrix $\Phi$ based on $({\mathbf{A}}_{t})$
${{\mathbf{T}}}_{t}=t^{\phi^{*}/2}g(\widetilde{A}_{t})^{-1/2}L^{11}(O^{11})^{\top}\,U\,V^{\top}\,{{\mathbf{A}}}_{t}=t^{\phi^{*}/2}g(\widetilde{A}_{t})^{-1/2}L^{11}(O^{11})^{\top}\,\widehat{{{\mathbf{A}}}}_{t}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},I_{N-1}\right)$
@@
In summary, whenever ${\mathbf{u}}_{0}={\mathbf{u}}_{1}$ we have under $H_{1}$ that ${{\mathbf{T}}}_{t}$ is still asymptotically normally distributed with covariance matrix $L^{11}(O^{11})^{\top}UV^{\top}{\mathcal{M}}_{det}^{11}VU^{\top}O^{11}(L^{11})^{\top}$, where ${\mathcal{M}}_{det}^{11}$ is referred to the eigen-structure of $\Phi_{1}$, while $L^{11}(O^{11})^{\top}UV^{\top}$ is referred to $\Phi_{0}$. Thus, the distance between the identity $I$ and the matrix $L^{11}(O^{11})^{\top}UV^{\top}{\mathcal{M}}_{det}^{11}VU^{\top}O^{11}(L^{11})^{\top}$ describes the relation between the asymptotic distribution of $\|{{\mathbf{T}}}_{t}\|^{2}$ under $H_{0}$ and the one under $H_{1}$, which determines the power of the test. For instance, note that $E[\|{{\mathbf{T}}}_{t}\|^{2}]=(N-1)$ under $H_{0}$, while $E[\|{{\mathbf{T}}}_{t}\|^{2}]$ is equal to the trace of $L^{11}(O^{11})^{\top}UV^{\top}{\mathcal{M}}_{det}^{11}VU^{\top}O^{11}(L^{11})^{\top}$ under $H_{1}$.
@@ B.4. Hypothesis tests on the matrix $\Phi$ based on $({\mathbf{B}}_{t})$
${{\mathbf{T}}}_{t}=t^{\phi^{*}/2}g(\phi^{*}\widetilde{B}_{t})^{-1/2}L^{33}(O^{33})^{\top}{\mathbf{B}}^{\prime}_{t}=t^{\phi^{*}/2}g(\phi^{*}\widetilde{B}_{t})^{-1/2}L^{33}(O^{33})^{\top}\,U\,V^{\top}\,{{\mathbf{B}}}_{t}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},I_{N-1}\right)$
@@ Appendix D Technical results
For the reader’s convenience, we here recall some results used in the above proofs. (For a synthetic review about the notion of stable convergence and its variants, see for instance the appendix of Aletti et al., (2020) and the references recalled therein.)
Theorem D.1 (A consequence of (Crimaldi,, 2009, Theorem A.1)).
@@
Lemma D.2 ((Aletti et al.,, 2024, Lemma S2.2)).
@@
Theorem D.3 (A consequence of (Crimaldi and Pratelli,, 2005, Proposition 3.1)).
@@
Recall that, according to the notation used here, the above condition 2) means that $\sup_{m_{0}\leq k\leq t}|s_{t}{\mathbf{T}}_{t,k+1}|=s_{t}\sup_{m_{0}\leq k\leq t}\sum_{i=1}^{d}|T_{t,k+1,i}|\leq Cy_{t}$ for a suitable deterministic constant $C$ and $t$ large enough, so that we trivially have the convergence in $L^{1}$ of $\sup_{m_{0}\leq k\leq t}|s_{t}{\mathbf{T}}_{t,k+1}|$ toward zero, as requested in (Crimaldi and Pratelli,, 2005, Proposition 3.1).
@@
Lemma D.4 (A consequence of (Berti et al.,, 2011, Lemma 1)).